# Hypergraph-MLP: Learning on Hypergraphs without Message Passing

### v2

@@ 4.2 Hypergraph-Smoothness-based Loss Function
Loss function via maximum likelihood estimation. Our goal is to create a loss function derived from the likelihood of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$, thus allowing the MLP-based model to utilize the hypergraph structural information by generating node embeddings that follow the smoothness prior mentioned above. With the distribution in Eq. (4) we have $p(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}},\mathbf{H})\!\propto\!e^{-\frac{1}{2}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}}$. Omitting the constant term, the negative log likelihood function of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ is:

### v3

@@ 4.2 Hypergraph-Smoothness-based Loss Function
Loss function via maximum likelihood estimation. Our goal is to derive a loss function from the likelihood of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$, thus allowing the MLP to utilize the hypergraph structural information by generating node embeddings that follow our smoothness prior. With Eq. (4) and setting $\mathbf{H}$ to obey a uniform distribution, we have $p(\mathbf{H}|\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})\!\propto\!e^{-\frac{1}{2}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]\mathbf{L}_{\mathbf{H}}[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}}$. Omitting the constant term, the negative log likelihood function of $[\mathbf{Z}_{\mathcal{V}}^{T},\mathbf{Z}_{\mathcal{E}}^{T}]^{T}$ is: